# A Comprehensive Survey of Deep Transfer Learning for Anomaly Detection in Industrial Time Series: Methods, Applications, and Directions

### v1

@@ 2.2 Deep transfer learning approaches
Since deep neural networks (DNNs) can learn useful feature representations from large amounts of data through back-propagation (Bengio et al., 2013), they have been widely adopted for tackling complex problems, which involve large-scale and high-dimensional data, also in practice (Schmidhuber, 2015; Lukic et al., 2016; Stadelmann et al., 2018; Schmidhuber, 2022). Deep transfer learning methods implement transfer learning principles within DNN and, among other things, enable deep learning based analysis pipelines to be applied to new datasets. On a high level, deep transfer learning approaches can be divided into data-driven and model-driven ones. Data-driven approaches focus on transferring knowledge by transforming and adjusting the data instances. Model-driven approaches leverage DNNs to develop domain-invariant features by reducing the feature discrepancy between source and target domain data and then transferring generalized knowledge to new tasks. Following the taxonomy of (Tan et al., 2018; Pan and Yang, 2010; Shi et al., 2022), we divide deep transfer learning approaches further into $4$ categories: instance transfer, parameter transfer, mapping transfer, and domain-adversarial transfer, as illustrated in Table 1. Instance transfer and mapping transfer are data-driven approaches, parameter transfer is a model-driven approach, while domain-adversarial transfer is a combination of both.
@@
Augmenting target data by transforming data instance from the source domain
He et al., 2022; Amirian et al., 2021; Wang et al., 2018
@@
Transfer learned parameters of a pre-trained model from source domain and adapting the model for target domain
Bommasani et al.,2022; Devlin et al.,2019; Vaswani et al.,2017; Dou et al.,2022; Alexandr et al.,2021; Yosinski et al.,2014; Guo et al.,2019; Sager et al.,2022; Brown et al.,2020
@@ 2.2.1 Instance transfer
The intuition of instance transfer is that although source and target domains differ, it is still possible to transform and reuse source data together with a few labeled target samples. A typical approach is to re-create some labeled data from the source domain. He et al. (2022) propose an instance-based deep transfer learning model with attention mechanism to predict stock movement. They first create new samples from the source dataset that are similar to the target samples by using attention weights, and then train on the created samples and target training samples for prediction tasks. Amirian et al. (2021) introduce an innovative instance transfer method for domain adaptation. They propose an effective auto-encoder model with a pseudo-label classifier to reconstruct new data instances that obtain general features across different datasets for medical image analysis. Taking another avenue, Wang et al. (2018) exclude the source data that have a bad impact on training target data. Specifically, they choose a pre-trained model from a source domain, estimate the impact of all training samples in the target domain, and remove samples that lower the performance of the model.
@@ 2.2.2 Parameter transfer
Parameter transfer adapts learned parameters of a pre-trained model to a new model. This assumes that DNNs can get similar feature representations from similar domains. Thus, through transferring parts of the DNN layers together with pre-trained parameters and/or hyperparameters, the pre-trained model is used as a base model to further train on target domain data and solve different learning tasks. Particularly, parameter transfer has gained popularity in computer vision and natural language processing, where large models are pre-trained on large datasets (Bommasani et al., 2022). In natural language processing, BERT (Devlin et al., 2019) and GPT-3 (Brown et al., 2020) that are based on the transformer architecture (Vaswani et al., 2017) can be fine-tuned for a variety of natural language processing tasks, including content generation (Dou et al., 2022), language translation (Sun et al., 2021), question answering (Glass et al., 2020), and summarization (Alexandr et al., 2021). Yosinski et al. (2014) investigated the general transferability of DNNs. Experiment results show that transferring features from source to target domain leads to improved generalization in networks compared to those trained solely on the target dataset. Unlike the typical way of fine-tuning a pre-trained model, Guo et al. (2019) propose the adaptive fine-tuning approach SpotTune to find the optimal fine-tuning strategy for the target task. Specifically, a policy network is used to make routing decisions on whether to pass the target instance through the pre-trained model. The results show SpotTune is effective in most cases by using a hybrid of parameter and instance transfer. Sager et al. (2022) propose an unsupervised domain adaptation for vertebrae detection in 3D CT volumes by transferring knowledge across domains during each batch of the training process.
@@ 2.2.3 Mapping transfer
Mapping transfer refers to learning a related feature representation for the target domain by feature transformation, which includes feature alignment, feature mapping, and feature encoding (Zhuang et al., 2021). The goal is to reduce feature discrepancies between source and target domains by minimizing the distance between the distribution of mapped features in the latent space. There are various criteria to measure the distribution difference, including Wasserstein distance (Shen et al., 2018), Kullback-Leibler Divergence (Dai et al., 2007), etc. Among them, Maximum Mean Discrepancy (MMD) (Tzeng et al., 2014) is most frequently adopted in mapping transfer from the surveyed papers. The MMD is calculated as the difference between the mean embeddings of the samples in a reproducing kernel Hilbert space associated with a chosen kernel function. Added to the target loss function, it measures the difference between two probability distributions and serves as a powerful tool for comparing the similarity of complex, high-dimensional datasets using a wide variety of kernel functions.
Some previous work focusing on transferred feature extraction/dimensionality reduction using MMD has been done. Long et al. (2017) base their Joint Adaptation Network on MMD, in which the joint distributions of multiple domain-specific layers across domains are aligned. In addition, an adversarial training version was adopted to make distributions of the source and target domains more distinguishable. Similarly, Long et al. (2015) adopted multi-layer adaptation and proposed Deep Adaptation Networks (DAN). In DAN models, the first three convolutional layers are used to extract general features. For the last three layers, multi-kernel MMD is used to bridge the cross-domain discrepancy and learn transferable features. Zhang et al. (2015) also based on MMD and proposed a Deep Transfer Network in which two types of layers are used to obtain domain invariant features across domains. The shared feature extraction layers learn a shared feature subspace between the source and the target samples, and the discrimination layer is then used to match conditional distributions by classifier transduction. Venkateswara et al. (2017) proposed Deep Adaptation Hash network, which is fine-tuned from the VGG-F (Chatfield et al., 2014) network. Multi-kernel MMD loss trains the Deep Adaptation Hash to learn feature representations that align the source and target domains.
@@ 2.2.4 Domain-adversarial transfer
Inspired by Generative Adversarial Networks (GANs) (Goodfellow et al., 2014; Schmidhuber, 2020), the goal here is to extract a transferable feature representation that is indiscriminative between source and target domain through adversarial training. Adversarial transfer mainly focuses on domain adaptation problems. Tzeng et al. (2015) adopt a domain confusion loss across the source and target domains to learn a domain invariant representation. Ganin et al. (2017) propose a new domain adaptation architecture by adding a domain classifier after feature extraction layers. A gradient reversal layer is used to ensure the similarity of the feature distributions over source and target domains. Ajakan et al. (2015) propose a domain adversarial DNN in which a domain regressor is applied to learn a domain invariant feature representation. Tzeng et al. (2017) use an unsupervised domain adaptation method that combines adversarial learning with discriminative feature learning.
3 Time series analysis for industrial processes
Time series analysis encompasses statistical techniques to analyze and interpret sequential temporal data. In the context of industrial processes, time series analysis plays a crucial role in automating monitoring and controlling the efficiency, quality, and performance of these processes. Specifically, the analysis of time series data can be used for anomaly detection, forecasting, process control, performance assessment, and maintenance scheduling to increase the efficiency of the process.
Anomaly detection
According to Hawkins (1980), an outlier is as an observation that deviates significantly from other observations in a way that it is likely that it was generated by a different mechanism. In this survey, we focus on time series data collected from machine sensor readings in the context of industrial applications, either univariate (only one variable is recorded over time) or multivariate (several simultaneously recorded measurements). Time series anomalies might occur for various reasons, including internal factors (e.g., temporary sensor error, machinery malfunction) and external factors (e.g. human error, ambient temperature). They can be divided into three categories (Choi et al., 2021; Chalapathy and Chawla, 2019): point anomalies, contextual anomalies, and collective anomalies. Point anomalies are isolated samples that deviate significantly from the normal behavior of that time series, which can be seen on the left of Fig. 3, e.g., a sudden spike in a pressure reading from a manufacturing machine sensor. These point anomalies can be caused by temporal sensor error, human error, or abnormal machinery operations. Contextual anomalies represent data points that deviate from normal ones only in their current context, and an example can be seen in the middle of Fig. 3. Collective anomalies are a set of data points that in their entirety (but not individually) are abnormal with respect to the entire time series, as shown on the right of Fig. 3.
Figure 3: Three time series anomaly types. Gray lines represent recorded time series signals, and dashed green lines are a priori set thresholds of normal operations. The red dots and the red line indicate anomalies. Point anomalies are single values that fall outside of a pre-set range (left panel). Contextual anomalies are samples that deviate from current context (middle panel). Collective anomalies are defined as a series of data points that all fall within the range of operation but jointly are not expected (right panel).
Challenges regarding detecting time series anomalies persist due to two specific properties: First, the complexity of time series data. As the automation level of industrial processes and the complexity of industrial systems increases, univariate time series data become insufficient and inefficient in representing any industrial process in its entirety. Hence, more sensors are installed to monitor the whole process, making it necessary in turn to detect anomalies from multivariate time series, which poses particular challenges since it requires consideration of temporal dependencies and relationships between variables and modalities. Second, the dynamic variability of industrial processes can pose difficulties in detecting anomalies due to fluctuations in the process caused by varying input or environmental conditions such as material, temperature, pressure, and humidity, which lead to domain shifts.
Process automation
@@ 3 Time series analysis for industrial processes
The tasks described above are combined to automate industrial processes. For example, after the detection of an anomaly, another model that captures the relationship between time course and different failure modes or drifts may be exploited for predictive maintenance. For example, in injection molding process monitoring, anomaly detection models are used to analyze recorded sensor data from injection molding machines to detect bad parts and identify the root cause of anomalies (Tercan et al., 2018). There are two basic ways to detect anomalies: for supervised anomaly detection, labels (normal/abnormal) are needed per time series to build a binary classifier (Görnitz et al., 2013). For unsupervised anomaly detection, an anomaly score or confidence value that is conditioned purely on normal data can be used to differentiate abnormal from normal instances (Zhang et al., 2023; Audibert et al., 2020; Stadelmann et al., 2019).

### v2

@@ 2.3 Deep transfer learning approaches
Since deep neural networks (DNNs) can learn useful feature representations from large amounts of data through back-propagation [24], they have been widely adopted for tackling complex problems in practice [38, 39, 40, 41], which involve large-scale and high-dimensional data. Deep transfer learning methods implement transfer learning principles within DNN and, among other things, enable deep learning based analysis pipelines to be applied to new datasets.
Based on the transferring techniques in the surveyed literature, we access how knowledge is shared across domains and help increase the performance in the target task or domain. we divide deep transfer learning approaches further into $4$ categories: instance transfer, parameter transfer, mapping transfer, and domain-adversarial transfer, as illustrated in Table 1. Furthermore, instance transfer, mapping transfer, and domain-adversarial transfer can be described as data-driven approaches. They focus on transferring knowledge by leveraging a large amount of data. It usually involves transforming and adjusting the data instances or manipulating data from different domains by feature alignment, feature mapping, etc. On the other hand, parameter transfer is a model-driven approach, which places more emphasis on understanding the underlying structure and dynamics of the data. It usually involves transferring the parameters of pre-trained model from source domain to target domain.
@@
Augmenting target data by transforming data instance from the source domain to the target domain
[42, 43, 44]
@@
Transfering learned parameters of a pre-trained model from source domain and adapting the model for target domain
[45, 46, 47, 48, 49, 50, 17, 51, 52, 53]
@@ 2.3.1 Instance transfer
The intuition of instance transfer is that although source and target domains differ, it is still possible to transform and reuse source data together with a few labeled target samples. A typical approach is to re-create some labeled data from the source domain. For example, He et al. propose an instance-based deep transfer learning model with an attention mechanism to predict stock movement [42]. They first create new samples from the source dataset that are similar to the target samples by using an attention network and then train another network on the created samples and target training samples for prediction tasks. Since two networks are trained separately for different tasks, it needs further investigation to what extent the generated samples can contribute to the prediction task. Amirain et al. introduce an innovative instance transfer method for domain adaptation [43]. They propose an effective auto-encoder model with a pseudo-label classifier to reconstruct new data instances that obtain general features across different datasets for medical image analysis. Taking another avenue, Wang et al. exclude the source data that negatively impacts training target data. Specifically, they choose a pre-trained model from a source domain, estimate the impact of all training samples in the target domain, and remove samples that lower the model’s performance. Then, the optimized training data is used for fine-tuning. The experiments are conducted on large image datasets [44]. Instead of transferring the data, the approach excludes certain samples based on the pre-trained model’s predictions. Additional validation is required in industrial environments, especially when only a few data are available in some industrial settings.
@@ 2.3.2 Parameter transfer
Parameter transfer adapts the learned parameters of a pre-trained model to a new model. This assumes that DNNs can get similar feature representations from similar domains. Thus, through transferring parts of the DNN layers together with pre-trained parameters and/or hyperparameters, the pre-trained model is used as a base model to further train on target domain data and solve different learning tasks. Particularly, parameter transfer has gained popularity in computer vision and natural language processing, where large models are pre-trained on large datasets [45]. In natural language processing, for example, BERT [46] and GPT-3 [53] are based on the Transformer architecture [48] which can be fine-tuned for a variety of downstream tasks, including content generation [49], language translation [66], question answering [67], and summarization [50]. In computer vision, Yosinski et al. investigate the general transferability of CNNs in image recognition [30]. They analyze the transferring effect by fine-tuning or freezing a certain amount of layers in the networks. Experimental results show that transferring features from source to target domain improves network generalization compared to those trained solely on the target dataset. Additionally, they quantify the model performance by assessing how features at what layers transfer from one task to another. It is surprising to find that transferring a pre-trained network from any number of layers can produce a boost for fine-tuning on a new dataset. However, the experiments are only conducted on certain image datasets, and Tuggener et al. show [68] the limits of parameter transfer when the chosen architecture is overfitted on the particularities of certain large-scale datasets.
Unlike the typical way of fine-tuning a pre-trained model, Guo et al. propose an adaptive fine-tuning approach SpotTune to find the optimal fine-tuning strategy for the target task [51]. Specifically, a policy network is used to make routing decisions on whether to pass the target instance through the pre-trained model. The results show SpotTune is effective in most cases by using a hybrid of parameter and instance transfer. Sager et al. propose an unsupervised domain adaptation for vertebrae detection in 3D CT volumes by transferring knowledge across domains during the training process [52].
@@ 2.3.3 Mapping transfer
Mapping transfer refers to learning a related feature representation for the target domain by feature transformation, which includes feature alignment, feature mapping, and feature encoding [27]. The goal is to reduce feature discrepancies between source and target domains by minimizing the distance between the distribution of latent feature representation. There are various criteria to measure the distribution difference, including Wasserstein distance [69], Kullback-Leibler Divergence [70], etc. Among them, Maximum Mean Discrepancy (MMD) [55] is most frequently adopted in mapping transfer from the surveyed papers. The MMD is calculated as the difference between the mean embeddings of the samples in a reproducing kernel Hilbert space associated with a chosen kernel function. Added to the target loss function, it serves as a powerful tool for comparing the similarity of complex, high-dimensional datasets using a wide variety of kernel functions.
Previous work has focused on transferred feature extraction/dimensionality reduction using MMD. Wang et al. focus more on the subdomain of the same subcategory instead of the alignment of the global distribution between source and target domain [54]. Specifically, they first use the attention mechanism to extract discriminative features that are most related to the fault signal. Then, local MMD is applied to transfer knowledge to adjust the distribution of related subdomains under the same category. Long et al. propose their Joint Adaptation Network [56] based on MMD, in which the joint distributions of multiple domain-specific layers across domains are aligned. In addition, an adversarial training version was adopted to make distributions of the source and target domains more distinguishable. Similarly, Long et al. adopt multi-layer adaptation and proposed Deep Adaptation Networks (DAN) [57]. The first three convolutional layers are used in DAN models to extract general features. For the last three layers, multi-kernel MMD bridges the cross-domain discrepancy and learns transferable features. Zhang et al. propose a Deep Transfer Network in which two types of layers are used to obtain domain invariant features across domains by adding MMD loss. The shared feature extraction layers learn a shared feature subspace between the source and the target samples, and the discrimination layer is then used to match conditional distributions by classifier transduction [58]. Venkateswara et al. propose Deep Adaptation Hash network [59], which is fine-tuned from the VGG-F [71] network. Multi-kernel MMD loss is employed to train the Deep Adaptation Hash network to learn feature representations that align the source and target domains.
@@ 2.3.4 Domain-adversarial transfer
Inspired by Generative Adversarial Networks (GANs) [72, 73], the goal of domain-adversarial transfer is to extract a transferable feature representation that is indiscriminative between source and target domain through adversarial training. Adversarial transfer is primarily concerned with addressing domain adaptation problems.
@@
Soleimani and Nazerfard utilize the GANs framework to perform cross-subject transfer learning [60]. The generator is used to generate samples that are similar to the target data. Meanwhile, the discriminator distinguishes the fake samples from the target samples. The classifier is trained to discriminate the labeled source data and fake samples to learn generalized features invariant to source and target domains. It is important to note that in real-world applications, training GANs can be unstable due to mode collapse, especially in the case when the source data and target data are unbalanced, and the generator may fail to generate fake samples that can confuse the discriminator. Tzeng et al. adopt a domain confusion loss across the source and target domains to learn a domain invariant representation [61]. Ganin et al. propose a new domain adaptation architecture by adding a domain classifier after feature extraction layers [63]. A gradient reversal layer is used to ensure the similarity of the feature distributions over source and target domains. Similarly, Ozyurt et al. develop a novel framework for unsupervised domain adaptation of time series data by using contrastive learning and domain-adversarial transfer learning [62]. A domain classification loss is applied to extract domain invariant features. The drawback is that the experiments are designed in a way that the source and target data sizes are similar, whereas in practice, the target data is usually much fewer than the source data. Ajakan et al. propose a domain adversarial DNN in which a domain regressor is applied to learn a domain invariant feature representation [64]. Tzeng et al. use an unsupervised domain adaptation method that combines adversarial learning with discriminative feature learning [65].
2.4 Related learning paradigms
Besides the dedicated transfer learning approaches discussed above, there are methods that represent alternative ways to solve tasks across domains or are complementary to the native transfer learning methods.
  • Multi-task learning is a machine learning technique where a single model is trained on multiple tasks simultaneously. The idea is to improve the performance of the model by learning a shared representation that captures the features between all tasks. Because the network learns to solve multiple tasks, it may generalize better to new data and tasks.
  • Continuous learning [74] is a learning process where the model continuously learns new tasks from previous tasks over time without forgetting how to solve previous tasks. To some extent, continuous learning can be seen as a sequential transfer learning process, with the constraint to preserve the performance of the previous tasks, which leads to an accumulation of knowledge over time.
  • Few-shot learning [75] is a type of machine learning where a model can learn and perform well on a new task with only a limited number of labeled samples. In extreme cases, the model can learn with one label [76] and without any label [77]. Whereas, transfer learning usually involves reusing the model from relevant tasks and continuing training on the target dataset.
  • Domain generalization[78, 79] focuses on developing a generalized model from one or multiple distinct domains to detect unseen target domain data. The main goal is to overcome the domain shift problem. Domain generalization and transfer learning are both applied to transfer knowledge from source domain to target domain. The major difference between transfer learning and domain generalization lies in the utilization of target domain data. Transfer learning leverages knowledge from source domain and target domain. In contrast, domain generalization solely learns from source domain, without access to the target data.
  • Meta-learning [80, 81] is known as “learning to learn”. For meta-learning, models are trained on a different set of tasks instead of a set of data in the traditional machine learning setting. In this sense, meta-learning can be seen as a form of transfer learning because it involves transferring knowledge from task to task.
  • Knowledge distillation [82] effectively learns a small model trained to mimic the behavior of a larger, more complex model. The knowledge learned by the larger model can be transferred to the smaller model, which can then be used for the target task.
  • Self-supervised learning [83, 84] involves training a model to predict some aspect of the input data without any external supervision. The learned representations can be used for various downstream tasks, including those that involve transferring knowledge from one domain to another.
3 Time series anomaly detection in industry
@@
Time series anomaly detection encompasses statistical techniques to analyze and interpret sequential temporal data. In the context of industrial processes, time series anomaly detection plays a crucial role in automating monitoring, effectively scheduling maintenance, and controlling the efficiency, quality, and performance of these processes. For example, after the detection of an anomaly, another model that captures the relationship between time course and different failure modes or drifts may be exploited for predictive maintenance. For example, in injection molding process monitoring, anomaly detection models are used to analyze recorded sensor data from injection molding machines to detect bad parts and identify the root cause of anomalies [85]. There are two basic ways to detect anomalies: for supervised anomaly detection, labels (normal/abnormal) are needed per time series to build a binary classifier [86]. For unsupervised anomaly detection, an anomaly score or confidence value that is conditioned purely on normal data can be used to differentiate abnormal from normal instances [87][88].
3.1 Anomaly types
According to the literature [89], an outlier is an observation that deviates significantly from other observations in a way that it is likely that it was generated by a different mechanism. In this survey, we focus on time series data collected from machine sensor readings in the context of industrial applications, either univariate (only one variable is recorded over time) or multivariate (several simultaneously recorded measurements). Time series anomalies might occur for various reasons, including internal factors (e.g., temporary sensor error, machinery malfunction) and external factors (e.g. human error, ambient temperature). They can be divided into three categories [32], [33]: point anomalies, contextual anomalies, and collective anomalies. Point anomalies are isolated samples that deviate significantly from the normal behavior of that time series, which can be seen on the left of Fig. 5, e.g., a sudden spike in a pressure reading from a manufacturing machine sensor. These point anomalies can be caused by temporal sensor error, human error, or abnormal machinery operations. Contextual anomalies represent data points that deviate from normal ones only in their current context, and an example can be seen in the middle of Fig. 5. Collective anomalies are a set of data points that in their entirety (but not individually) are abnormal with respect to the entire time series, as shown on the right of Fig. 5.
Figure 5: Three time series anomaly types. Gray lines represent recorded time series signals, and dashed green lines are a priori set thresholds of normal operations. The red dots and the red line indicate anomalies. Point anomalies are single values that fall outside of a pre-set range (left panel). Contextual anomalies are samples that deviate from the current context (middle panel). Collective anomalies are defined as a series of data points that all fall within the range of operation but jointly are not expected (right panel).
3.2 Challenges
Challenges regarding detecting time series anomalies persist due to two specific properties: (1) The complexity of time series data. As the automation level of industrial processes and the complexity of industrial systems increases, univariate time series data become insufficient and inefficient in representing any industrial process in its entirety. Hence, more sensors are installed to monitor the whole process, making it necessary to detect anomalies from multivariate time series, which poses particular challenges since it requires consideration of temporal dependencies and relationships between variables and modalities. Many researchers work on discovering generalized patterns from spatial and temporal correlated multivariate time series data. Zhang et al. propose a Deep Convolutional Autoencoding Memory network [87], where they build an autoencoder to capture spatial dependency of multi-variant data using MMD to distinguish noisy, normal and abnormal data. Zhu et al. propose an interpretable model agnostic multivariate time-series anomaly detection method for applications of cyber physical systems [90]. The new method considers both the temporal and feature dimensions through an adaptive mask based series saliency module to produce accurate anomaly detection results and reasonable interpretations in the form of a mask matrix. (2) The dynamic variability in industrial processes. Industrial processes often have high dynamic variability and can be affected by a wide range of conditions, such as changes in temperature, pressure, and humidity. These conditions can cause fluctuations in the process outputs, which leads to data shift and domain shift. This can make it challenging to detect anomalies and maintain control over the industrial process.
3.3 Anomaly detection methods
Time series anomaly detection has been investigated for decades, and various types of methods have been proposed [91]. This paper exclusively discusses the time series anomaly detection techniques using deep learning, leveraging its robust representation learning capabilities. Current deep learning methods can be mainly divided into reconstruction-based, forecasting-based, and other methods. Fig. 6 illustrates the two main methods. In deep reconstruction-based anomaly detection, the reconstructed sequence is used to compare with the actual sequence. Differently, in deep forecasting-based anomaly detection, only the forecasted sequence is used to assess the similarity to the ground truth.
Figure 6: Illustration of deep learning-based anomaly detection (top row) with reconstruction-based (center row) and forecasting-based (bottom row) anomaly detection in time series. The first column represents two time series. The second column shows the reconstructed (top) and forecasted (bottom) time series. The third column shows the difference between reconstructed/forecasted time series. Deviations from the reconstructed or forecasted time series are indicative of an anomaly. In deep reconstruction-based anomaly detection, the entire sequence is reconstructed in a decoder-encoder architecture, and the reconstructed sequence is used to compare with the actual sequence. In deep forecasting-based anomaly detection, the end of a sequence is predicted using the start of the sequence and only the forecasted sequence is used to assess the similarity to the ground truth. In this example, the red time series has a likely anomaly at about 1.5 seconds. (Best viewed in color.)
Reconstruction-based methods
Reconstruction-based methods aim to learn the data distribution of the normal time series and differentiate the abnormalities from the normal ones by computing the reconstruction errors. Audibert et al. propose a fast and stable method – Unsupervised anomaly detection for multivariate time series [92], based on adversely trained autoencoders. The encoder-decoder architecture within an adversarial training framework combines the advantages of autoencoders and adversarial training while compensating for the limitations of each technique. After training two autoencoders, the anomaly score is defined by balancing the reconstructed errors from the two autoencoders with two hyperparameters. However, the challenge arises in selecting these two hyperparameters of the anomaly score when the testing dataset is unavailable.
Malhotra et al. also formulate an anomaly score based on reconstruction error [93]. They first train the LSTM encoder-decoder model to reconstruct the normal time series. Subsequently, they leverage the reconstruction errors to calculate the probability by using Maximum Likelihood Estimation to detect a specific point within a time series as an anomaly. They set a window to detect anomalies. The window will be labeled as anomalous if the probability exceeds a threshold. Similarly, Wei et al. also propose an LSTM-based encoder-decoder model to detect multivariant time series sequences based on the reconstruction error [94]. The major difference is the anomaly detection criteria. They assume the reconstruction error of train/test data follows the normal distribution and detect anomalies by using the 2-sigma rule of the normal distribution as a threshold. Zeng et al. propose an adversarial transformer structure to detect multivariate time series anomalies effectively [95]. Here, two-stage adversarial training is applied for the transformer. In the first stage, two transformers are trained by minimizing the reconstruction error to capture the temporal trends in the time series. In the second stage, the reconstruction error serves as prior knowledge in the adversarial training process, enabling the model to distinguish anomalies from normal time series. Then, an anomaly score is defined by combining the anomaly probability and reconstruction error. Again, a threshold has been chosen to differentiate anomalies from normal ones.
GANs, as effective unsupervised learning methods, have been used in time series anomaly detection. Anomaly detection methods based on GANs focus on extracting features by adversarial training on normal samples. Consequently, features from the abnormal samples diverge from those of the normal ones, reflecting in reconstruction error and discrimination value. Li et al. use LSTM-RNN as a base model for building generator and discriminator in GAN[96]. The proposed framework considers multiple variables to capture the temporal correlation of multi-time series distributions. Additionally, they proposed a novel anomaly score, which can detect anomalies through discrimination and reconstruction. More specifically, the score is a combination of the reconstruction difference between generated data and original data and the discrimination results from the discriminator. Similarly, Niu et al. and Bashar et al. both propose an LSTM-based VAE-GAN for time series anomaly detection, where LSTM networks are used as the generator, and discriminator [97, 98]. When it comes to anomaly scores, setting an optimal threshold is usually a critical step. However, using a small portion of the test set to decide the optimal threshold may not be practical in real-world scenarios [97]. Additionally, it is important to note that the method has been only tested for point anomaly detection, further investigation is required when they are applied to detect other anomaly types.
Forecasting-based methods
Forecasting-based methods predict the value of the following timestamps and predict temporal anomalies according to the prediction error. Kim et al. propose a forecasting-based unsupervised time-series anomaly detection method using transformer architecture [99]. The idea is to train a transformer-like model by forecasting a fixed-length time series based on the previous timestamps. The trained model is used to predict time series with an anomaly score such that an instance where the anomaly score is larger than a static threshold is defined as an anomaly. A dynamic thresholding technique is also mentioned but not explicitly discussed in the paper.
Deng and Hooi propose a novel attention-based graph neural network approach [100] that learns a graph of dependence relationships between multi-variant time series signals by forecasting the behavior based on past time series. Then, a graph deviation scoring is defined for each sensor to detect and explain anomalies. Tang et al. propose an interpretable multivariate time series anomaly detection method based on graph neural networks and gated recurrent units [101]. The feature representation is learned through forecasting the future time series segment. An abnormal score is set for each time series to detect anomalies. The feature embedding is then used for 2D visualization through t-SNE plots to interpret the clusters within time series from different sensors.
Other methods
Ding et al. propose a joint network to integrate the advantages of reconstruction and forecasting/prediction [102]. First, they propose a multimodal graph attention network to tackle the spatial-temporal dependencies for multimodal time series. Further, they optimize the reconstruction and prediction modules simultaneously to predict anomalies. Himeur et al. take advantage of annotated data and directly use a DNN as a classifier to classify normal and abnormal energy consumption types [103]. The enormous imbalance of real anomaly patterns is one concern in the approach. Thus, a normalized technique of power consumption data is applied to deal with this problem. The normalized data represent the difference in power consumption rates of each current time sample and the previous one. It can provide information on how fast the consumption reacts to the time evolution. However, any further evaluation of this technique is not discussed, and it is still an open problem regarding anomaly detection for other datasets. Yang et al. propose a contrastive learning structure with dual attention to learn a permutation invariant representation of the data with superior discrimination characteristics between normal points and anomalies [104]. Unlike most reconstruction-based models, their model is a self-supervised framework based on representation learning. The new method achieves state-of-the-art comparable performance on six multivariate and one univariate time series anomaly detection benchmark datasets. However, the extensive framework with two multi-head-attention blocks may be prone to overfitting. This concern is amplified by the absence of training details, leaving only evaluation details disclosed.
In principle, these anomaly detection approaches are applicable to all types of anomalies. Reconstruction-based methods are typically applied to the entire or a portion of the time series. Long-time series are commonly segmented into subsequences using a predefined sliding window. In the case of detecting context/collective anomalies, the reconstructed loss of the time series sequence is evaluated within the predefined sliding window, if the reconstruction loss is larger than an acceptable threshold, then that time series sequence is classified as an anomaly. In the case of point anomalies, reconstruction is performed at each single time stamp, akin to a regression problem, and then the reconstruction loss of each single timestamp is evaluated to determine whether the single timestamp is anomalous or not. It is also applied to forecasting-based anomaly detection methods, instead of computing reconstruction error, forecasting-based methods predict the value in the next time stamp for point anomalies or the next time series sequence for context/collective anomalies. The anomalies will be detected based on the deviation between the predicted value and the normal value. Other anomaly detection approaches usually combine the reconstruction-based and foresting-based methods.
To sum up, these methods are applicable to each type of anomaly. However, the effectiveness of these anomaly detection approaches may vary depending on the anomaly detection tasks at hand, which are characterized by the granularity at which the time series data is observed and analyzed.